# The mean solar butterfly diagram and poloidal field generation rate at the surface of the Sun

### v1

@@ 5 Conclusion
In this paper we have shown that one must be careful when comparing butterfly diagrams produced by simple mean-field models with observed ones. Because of imperfect scale separation, the use of the azimuthal average does not result in an observed butterfly diagram that can be understood in the mean-field sense. The observed butterfly diagram must for this be cycle-averaged, a procedure made possible by the universality of the equatorward sunspot belt migration.
@@
By making use of the average toroidal field butterfly diagram and Joy’s law, we found the $\alpha$-effect-like BL source term is qualitatively consistent in shape and amplitude with the source term as inferred from the SFT model and the radial field butterfly diagram. This consistency is further increased when we assume the toroidal field emergence velocity to be dependent on the toroidal field strength itself.
@@ Acknowledgements.
This work was carried out when SC was a member of the International Max Planck Research School for Solar System Science at the University of Göttingen. The authors acknowledge partial support from ERC Synergy grant WHOLE SUN 810218.

### v2

@@ 6 Conclusions
In this paper we have shown that one must be careful when comparing butterfly diagrams produced by simple mean-field models with observed ones. Because of imperfect scale separation, the use of the azimuthal average does not result in an observed butterfly diagram that can be understood in the mean-field sense. The observed butterfly diagram must for this purpose be cycle-averaged, a procedure made possible by the universality of the equatorward sunspot belt migration. From this standard law we can determine a reference time when all cycles are in phase with each other and can thus be meaningfully averaged. To ‘increase’ the small number of cycles for which magnetogram observations exist by a factor of two, we symmetrized both hemispheres together.
@@
By making use of the average toroidal field time-latitude diagram and Joy’s law, we found that the $\alpha$-effect-like BL source term is qualitatively consistent in shape and amplitude with the source term as inferred from the SFT model and the butterfly diagram. This consistency is further increased when we assume the toroidal field emergence velocity to be dependent on the toroidal field strength itself.
@@ Acknowledgements.
The authors wish to thank the anonymous referee for comments that helped improve the overall quality of this paper. This work was carried out when SC was a member of the International Max Planck Research School for Solar System Science at the University of Göttingen. The authors acknowledge partial support from ERC Synergy grant WHOLE SUN 810218.